OpenStudy (anonymous):

Find the derivative of V= 1000 ( 1- t/60 ) ^2 using limits

OpenStudy (bahrom7893):

shappy, go for it!

OpenStudy (anonymous):

haha

OpenStudy (bahrom7893):

NotSObright's got this!

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{1000(1-\frac{t+h}{60})^2-1000(1-\frac{t}{60})^2}{h}\]

OpenStudy (anonymous):

This is righthand derivative

OpenStudy (anonymous):

yea i got this part too.. but then I messed up the other steps

OpenStudy (anonymous):

Left hand derivative is\[\lim_{h \rightarrow 0}\frac{1000(1-t/60)^2-1000(1-\frac{t-h}{60})^2}{}\]

OpenStudy (anonymous):

I am using the right hand derivative i expanded it..but then got 0 as my final answer so dont know what is wrong